# Comfortability of quantum walks on embedded graphs on surfaces

### v1

@@ 4.4 Extension to an infinite system
For the blow-up graph $G(\rho,\tau)$, let us choose a subset of $A_{is}$ by $\delta A_{is}$ as the boundary to the “outside”. We deform this blow-up graph to an infinite graph by the following procedure.
@@
    Every $e\in A(\rho,\tau)\setminus\delta A_{is}$ is left as it is without any transformations.
@@
    Each $e\in\delta A_{is}$ is divided by two arcs by replacing $e$ with two new arcs $\xi_{out}$ and $\xi_{in}$ satisfying with $o(e)=o(\xi_{out})$, $t(\xi_{out})=o(\xi_{in})$ and $t(\xi_{in})=t(e)$.
@@
    Semi-infinite path $\mathbb{P}_{e}$ with the root vertex $o(\mathbb{P}_{e})$ is joined by identifying $o(\mathbb{P}_{e})$ with $t(\xi_{out})=o(\xi_{in})$. The pair of $(\xi_{out},\xi_{in})$ is called the quay of $e\in\delta A_{is}$. The sets of such as $\xi_{out}$ and $\xi_{in}$ are denoted by $\delta A_{qy}^{-}$ and $\delta A_{qy}^{+}$, respectively.
@@
The set of arcs of tails is denoted by $A_{tl}$. The subset of $A_{tl}$ called the pier is denoted by
$\delta A_{pr}^{+}=\{e\in A_{tl}\;|\;t(e)\in\delta X\},\;\delta A_{pr}^{-}=\{e\in A_{tl}\;|\;o(e)\in\delta X\}.$
Note that for each island having a boundary vertex, the local rotation in $\tilde{G}(\rho,\tau)$ is “expanded” by the insertion of the tail. Let us use the same notation for the new rotation as $\rho$ to reduce the number of notations. Let $\tilde{A}_{is}\subset{\tilde{A}}^{BU}$ be the set of the islands of $\tilde{G}(\rho,\tau)$ following the rotation $\rho$. Let $\tilde{A}_{br}$ be the union of $A_{br}$ and $\delta A_{pr}$. The designation of the twist to $A_{tl}$ is assigned by 00.
@@
The time evolution operator $\tilde{W}$ on $\mathbb{C}^{\tilde{A}(\rho,\tau)}$ is given as follows, which is essentially the same as Definition 4.1 except the boundaries and the tails but the setting of the initial state is crucial to obtain the stationary state:
@@
Let $\Psi_{t}$ be the $t$-th iteration of $\tilde{W}$, such that $\Psi_{t}=\tilde{W}\Psi_{t-1}$ $t=1,2,\dots$ with the initial state $\Psi_{0}$ defined in (4.17).
The time evolution $\tilde{W}$.
@@
    For $\boldsymbol{e}\in A_{tl}\setminus\delta A_{pr}^{-}$: The arcs of a tail is put by $e_{0},e_{1},\dots$ and $\bar{e}_{0},\bar{e}_{1},\dots$ with $t(e_{0})\in\delta X$ $t(e_{j+1})=o(e_{j})$ $j=0,1,2,\dots$. The time evolution on the tail is free, that is,
@@
    For $\boldsymbol{e}\in\tilde{A}_{is}\cup A_{br}\cup\delta A_{pr}^{-}$: For any island arc $\boldsymbol{e}_{is}\in\tilde{A}_{is}$, “bridge or pier” arc $\boldsymbol{e}_{br}\in A_{br}\cup\delta A_{pr}^{-}$,
@@
Let each tail is labeled by each element of $\delta A_{pr}^{+}$. Prepare the sequence of complex values $(\alpha_{e})_{e\in\delta A_{pr}^{+}}$. Then we set the following uniformly bounded initial state on the tails:
$\Psi_{0}(e^{\prime})=\begin{cases}\alpha_{e}&\text{: $e^{\prime}\in A(\mathbb{P}_{e})$, ${\rm dist}(t(e^{\prime}),o(\mathbb{P}_{e}))>{\rm dist}(t(e^{\prime}),o(\mathbb{P}_{e}))$ ($e\in\delta A_{is}$),}\\ 0&\text{ otherwise.}\end{cases}$ (4.17)
Note that the initial state $\Psi_{0}$ is bounded but no longer square summable. Such a setting of initial state provides a constant inflow from the tails at very time step. On the other hand, the setting of the time evolution on tails implies an outflow from the interior. Under this setting, the convergence to a fixed point of this quantum walk is ensured in the following meaning:
@@ 5 Unitary equivalence
Choose a one vertex from $G$, say $x\in X$. Then let us reverse its rotation and switch all the edge types of the incident edges of $x$. Such a rotation system is denoted by $(G,\rho^{(x)},\tau^{(x)})$.
By Lemma 4.1, the resulting embeddings on a closed surface of the rotation systems of the double covering graph, $(G^{\tau},\rho\oplus\rho^{-1},\mathrm{id})$ and $(G^{\tau^{(x)}},\rho^{(x)}\oplus{\rho^{(x)}}^{-1},\mathrm{id})$, are isomorphic to each other. Under these isomorphic two embeddings, we take the blow up and choose the boundary island arcs $\delta A$ from $G(\rho,\tau)$ and its isomorphic boundary $\delta A^{\prime}$ from $G(\rho^{(x)},\tau^{(x)})$. The resulting blow up graphs are denoted by $\tilde{G}(\rho,\tau)$ and $\tilde{G}(\rho^{(x)},\tau^{(x)})$, respectively. Note that they are still isomorphic to each other: $\tilde{G}(\rho,\tau)\cong\tilde{G}(\rho^{(x)},\tau^{(x)})$ under the following bijection $\phi$ for the incident arcs of islands $x$ and $x^{\prime}$:
@@
    $e_{is}=(e,\tilde{\rho}(e))\in A_{is}(x,k)\subset A(\rho,\tau)$
    with $o(e)=(y,k^{\prime})$ and $o(\tilde{\rho}(e))=(z,k^{\prime\prime})$ in $G^{\tau}$
@@
    $\phi(e_{is})=(e^{\prime},\tilde{\rho}(e^{\prime}))\in A_{is}(x,k+1)\subset A(\rho,\tau)$
    with $o(e^{\prime})=(z,k^{\prime\prime}+1)$ and $t(\tilde{\rho}(e^{\prime}))=(y,k^{\prime}+1)$ in $G^{\tau}$.
@@
Let the rotation system $(G,\rho,\tau)$ and $(G,\rho^{\prime},\tau^{\prime})$ are isomorphic to each other, and let the boundaries of the induced blow-up graphs are also isomorphic to each other. Then the time evolution operators $W$ on $\tilde{G}(\rho,\tau)$ and $W^{\prime}$ on $\tilde{G}(\rho^{\prime},\tau^{\prime})$ are unitarily equivalent, that is, there is an unitary map $\mathcal{U}$ such that
@@
It is sufficient to consider the case for $\rho^{\prime}=\rho^{x}$ for arbitrary $x\in X$. By Lemma 4.1, “the reversing the rotation of $x$” and “the reversing all the incident edges of $x$” are reflected in the corresponding time evolution operator on $\tilde{G}(\rho,\tau)$:
  1. (1)
@@
Obviously, the operation (1) has the corresponding unitary map. This map is denoted by $\mathcal{U}_{\phi}$, that is,
@@
Then in the rest of the discussion, let us find the corresponding unitary map of the operation (2). To this end, for the blow-up graph $\tilde{G}(\rho,\tau)$, set
@@
$\theta(e)=\begin{cases}\pi&\text{: $e\in A_{br}$ with $\tau(e)=1$}\\ 0&\text{: otherwise}\end{cases}$ (5.18)
then $\tau(e)=e^{\boldsymbol{\rm i}\;\theta(e)}$ for any $e\in A_{br}$, where $\boldsymbol{\rm i}=\sqrt{-1}$. On the other hand if we set
$\theta(e)=\begin{cases}\pi&\text{: $e\in A_{br}$ with $\tau^{\prime}(e)=1$}\\ 0&\text{: otherwise}\end{cases}$ (5.19)
then $\tau^{\prime}(e)=e^{\boldsymbol{\rm i}\;\theta(e)}$ for any $e\in A_{br}$. For $\theta\in\mathcal{A}$, and arbitrary path in $G(\rho,\tau)$, $p=(a_{1},a_{2},\dots,a_{k})$ with $t(a_{j})=o(a_{j+1})$ $(j=1,\dots,k-1)$, we define
@@
A path $p=(a_{1},\dots,a_{r})$ with $a_{1}\in A_{br}$, $a_{r}\in A_{is}$ in $G(\rho,\tau)$ may be represented by
[…]
where $\xi_{j},\tilde{\rho}^{s}(\xi_{j})\in A_{is}$ and $e_{j}\in A_{br}$. Here if there is a back-tracking in $p$, a corresponding subsequence of a walk on an island disappears. The inverse path of $p$ is defined by
[…]
where $\tilde{\rho}:=\rho\oplus\rho^{-1}$. The inverse path of $p$ for the other cases and $p$ with back-trackings case are also defined in the same way.
@@
Set $\theta_{1},\theta_{2}\in\mathcal{A}$. If $\int_{c}\theta_{1}=\int_{c}\theta_{2}$ for any close path $c$ in $G(\rho,\tau)$, then $\int_{p}(\theta_{2}-\theta_{1})=\int_{p^{\prime}}(\theta_{2}-\theta_{1})$ for any path $p$ and $p^{\prime}$ with $o(p)=o(p^{\prime})$ and $t(p)=t(p^{\prime})$.
@@
It holds that $\int_{\bar{p}}\theta=-\int_{p}\theta$ for any path $p$ by the definition. Note that $p\cup\bar{p^{\prime}}$ is a closed cycle. Then we have
$0=\int_{p\cup\bar{p}}(\theta_{2}-\theta_{1})=\int_{p}(\theta_{2}-\theta_{1})+\int_{\bar{p^{\prime}}}(\theta_{2}-\theta_{1})=\int_{p}(\theta_{2}-\theta_{1})-\int_{p^{\prime}}(\theta_{2}-\theta_{1}),$
@@
Lemma 5.1 tells us that the value $\int_{p}(\theta_{2}-\theta_{1})$ is independent of routes from $o(p)$ to $t(p)$. Let us fix a vertex $x_{*}$. Since $\int_{p}(\theta_{2}-\theta_{1})$ for a path starting from $x_{*}$ is determined by $t(p)=x$, we set such a value by $\Delta(x)$, which is well-defined.
@@
for any $\epsilon\in A_{br}$, where $\xi^{\flat}=\mathrm{is}(\epsilon)$ and $\xi=\mathrm{is}^{\sharp}(\epsilon)$. Here $\theta_{1}$ is defined in (5.18). Now let us set $\theta_{2}$ by (5.19). The operation ()($\star$)( ⋆ ) keeps the parity of any cycles passing through the vertex $x$, because $\tau(e)+\tau(e^{\prime})$ with $t(e)=x=o(e)$ in $G$ is changed to
@@
Here $\Delta(x)$ with a fixed vertex $x_{*}\in\tilde{X}$ is well-defined by Lemma 5.1. Inserting them into (5.20), we have
@@
The second equality derives from $o(\xi)=o(\epsilon))$ which gives the commutativity of the second diagonal matrix and $C$ in first equality. Therefore introducing the unitary map $\mathcal{U}_{\Delta}$ by

### v2

@@ 4.4 Extension to an infinite system
For the blow-up graph $G(\rho,\tau)$, let us choose a subset of $A_{is}$ by $\delta A_{is}$ as the boundary to the “outside”. We deform this blow-up graph to an infinite graph by the following procedure. See Figure 11.
Definition 4.2 (The blow-up graph with tails).
The blow-up graph with tail induced by $G(\rho,\tau)$ is given as follows.
@@
    Each $\xi\in\delta A_{is}$ is divided by two arcs by replacing $\xi$ with two new arcs $\xi^{-}$ and $\xi^{+}$ satisfying with $o(\xi)=o(\xi^{-})$, $t(\xi^{-})=o(\xi^{+})$ and $t(\xi^{+})=t(\xi)$.
@@
    Semi-infinite path $\mathbb{P}_{e}$ with the root vertex $o(\mathbb{P}_{e})$ is joined by identifying $o(\mathbb{P}_{e})$ with $t(\xi^{-})=o(\xi^{+})$. The pair of $(\xi^{-},\xi^{+})$ is called the quay of $e\in\delta A_{is}$. The sets of such as all $\xi^{-}$ and all $\xi^{+}$ are denoted by $\delta A_{qy}^{-}$ and $\delta A_{qy}^{+}$, respectively.
@@
    Every $e\in A(\rho,\tau)\setminus\delta A_{is}$ is left as it is, without any transformations.
Figure 11: The quay and pier of the blow-up graph: The gray discs represent some islands of the blow-up graph. The semi-infinite path with the white vertices is the tail. The replaced arcs in the deformation process (1) with $\xi\in\delta A$ are called the quay arcs. The tail is inserted between the quay arcs $\xi^{-}$ and $\xi^{+}$. The arcs of of tails incident to the internal graph are called the pier arcs.
@@
The set of arcs of tails is denoted by $A_{tl}$. The subset of $A_{tl}$ called the pier is defined as
$\delta A_{pr}=\delta A_{pr}^{+}\sqcup\delta A_{pr}^{-},$
where
$\delta A_{pr}^{+}=\{t^{+}\in A_{tl}\;|\;t(t^{+})\in\delta X\},\;\delta A_{pr}^{-}=\{t^{-}\in A_{tl}\;|\;o(t^{-})\in\delta X\}.$
Note that for each island having a boundary vertex, the local rotation in $\tilde{G}(\rho,\tau)$ is naturally extended by the insertion of the tail. We use the same symbol $\rho$ for the new rotation. Let us introduce the island $\tilde{A}_{is}$ and the bridge $\tilde{A}_{br}$ in $\tilde{G}(\rho,\tau)$ as follows.
$\tilde{A}_{is}$ $=(A_{is}\setminus\delta A_{is})\sqcup\delta A_{qy}^{+}\sqcup\delta A_{qy}^{-},$
$\tilde{A}_{br}$ $=A_{br}\sqcup\delta A_{pr}^{+}\sqcup\delta A_{pr}^{-}.$
We assume $\tau(e)=0$ for any $e\in A_{tl}$.
@@
The time evolution operator $\tilde{W}$ on $\mathbb{C}^{\tilde{A}(\rho,\tau)}$ is given as follows, which is essentially the same as Definition 4.1 except the boundaries and the tails, but the setting of the initial state is crucial to obtain the stationary state:
@@
Let $\Psi_{t}$ be the $t$-th iteration of $\tilde{W}$, such that $\Psi_{t}=\tilde{W}\Psi_{t-1}$ $t=1,2,\dots$ Here the time eovlution operator $\tilde{W}$ and the initial state $\Psi_{0}$ are defined as follows.
The time evolution operator $\tilde{W}$.
@@
    For $\boldsymbol{e}\in A_{tl}\setminus\delta A_{pr}^{-}$: The arcs of a tail are put by $e_{0},e_{1},\dots$ and $\bar{e}_{0},\bar{e}_{1},\dots$ with $t(e_{0})\in\delta X$ $t(e_{j+1})=o(e_{j})$ $j=0,1,2,\dots$. The time evolution on the tail is free, that is,
@@
    For $\boldsymbol{e}\in\tilde{A}_{is}\cup A_{br}\cup\delta A_{pr}^{-}$: Let $a,b,c$ and $d$ are the same as in Definition 4.1. For any island arc $\boldsymbol{e}_{is}\in\tilde{A}_{is}$, and for any “bridge or pier” arc $\boldsymbol{e}_{br}\in A_{br}\cup\delta A_{pr}^{-}$,
@@
Let each tail be labeled by each element of $\delta A_{pr}^{+}$. Prepare the sequence of complex values $(\alpha_{e})_{e\in\delta A_{pr}^{+}}$. Then we set the following uniformly bounded initial state on the tails:
$\Psi_{0}(e^{\prime})=\begin{cases}\alpha_{e}&\text{: $e^{\prime}\in A(\mathbb{P}_{e})$, ${\rm dist}(t(e^{\prime}),o(\mathbb{P}_{e}))>{\rm dist}(t(e^{\prime}),o(\mathbb{P}_{e}))$ ($e\in\delta A_{is}^{+}$),}\\ 0&\text{ otherwise.}\end{cases}$ (4.20)
Note that the initial state $\Psi_{0}$ is bounded but no longer square summable. Such a setting of initial state provides a constant inflow from the tails at every time step. On the other hand, the setting of the time evolution on the tails implies that a walk never returns to the interior once it goes out to the outside, which can be regarded as an outflow from the interior. In this setting, the convergence to a fixed point of this quantum walk is ensured with the following meaning:
@@ 5 Unitary equivalence
Choose a one vertex from $G$, say $x\in X$. Then reverse its rotation and switch all the edge types of the incident edges of $x$. Such a rotation system is denoted by $(G,\rho^{(x)},\tau^{(x)})$.
By Lemma 4.1, the resulting embeddings on a closed surface of the rotation systems of the double covering graph, $(G^{\tau},\rho\oplus\rho^{-1},\mathrm{id})$ and $(G^{\tau},\rho^{(x)}\oplus{\rho^{(x)}}^{-1},\mathrm{id})$, are equivalent embedding to each other. Under these equivalent two embeddings, we take the blow-up and choose the boundary island arcs $\delta A$ from $G(\rho,\tau)$ and its isomorphic boundary $\delta A^{\prime}$ from $G(\rho^{(x)},\tau^{(x)})$. The resulting blow-up graphs are denoted by $\tilde{G}(\rho,\tau)$ and $\tilde{G}(\rho^{(x)},\tau^{(x)})$, respectively. Note that they are still isomorphic to each other: $\tilde{G}(\rho,\tau)\cong\tilde{G}(\rho^{(x)},\tau^{(x)})$ under the following bijection $\phi$ for the incident arcs of islands $x$ and $x^{\prime}$:
@@
    $e_{is}=(e,\rho(e))\in A_{is}((x,k))\subset A(\rho,\tau)$
    with $o(e)=(y,k^{\prime})$ and $o(\rho(e))=(z,k^{\prime\prime})$ in $G^{\tau}$
@@
    $\phi(e_{is})=(e^{\prime},\rho(e^{\prime}))\in A_{is}((x,k+1))\subset A(\rho,\tau)$
    with $o(e^{\prime})=(z,k^{\prime\prime}+1)$ and $o(\rho(e^{\prime}))=(y,k^{\prime}+1)$ in $G^{\tau}$.
@@
Let the rotation system $(G,\rho,\tau)$ and $(G,\rho^{\prime},\tau^{\prime})$ are equivalent embeddings to each other, and let the boundaries of the induced blow-up graphs are also equivalent to each other. Then the time evolution operators $W$ on $\tilde{G}(\rho,\tau)$ and $W^{\prime}$ on $\tilde{G}(\rho^{\prime},\tau^{\prime})$ are unitarily equivalent, that is, there is an unitary map $\mathcal{U}$ such that
@@
It is sufficient to consider the case for $\rho^{\prime}=\rho^{(x)}$ for a fixed arbitrary vertex $x\in X$. By Lemma 4.1, “the reversing the rotation of $x$” and “the reversing all the incident edges of $x$” are reflected in the corresponding time evolution operator on $\tilde{G}(\rho,\tau)$:
  1. Op(1)
@@ 2. Op(2)
Obviously, Op(1) has the corresponding unitary map. This map is denoted by $\mathcal{U}_{\phi}$, that is,
@@
Then in the rest of the discussion, let us find the corresponding unitary map of Op(2) following [13, Lemma 2.4]. To this end, for the blow-up graph $\tilde{G}(\rho,\tau)$, set
@@
$\theta_{1}(e)=\begin{cases}\pi&\text{: $e\in A_{br}$ with $\tau(e)=1$,}\\ 0&\text{: otherwise,}\end{cases}$ (5.21)
then $\theta_{1}\in\mathcal{A}$ and $e^{\boldsymbol{\rm i}\;\theta_{1}(e)}=(-1)^{\tau(e)}$ for any $e\in A_{br}$, where $\boldsymbol{\rm i}=\sqrt{-1}$. On the other hand if we set
$\theta_{2}(e)=\begin{cases}\pi&\text{: $e\in A_{br}$ with $\tau^{\prime}(e)=1$,}\\ 0&\text{: otherwise,}\end{cases}$ (5.22)
then $\theta_{2}\in\mathcal{A}$ and $e^{\boldsymbol{\rm i}\;\theta_{2}(e)}=(-1)^{\tau^{\prime}(e)}$ for any $e\in A_{br}$. For $\theta\in\mathcal{A}$, and an arbitrary path in $G(\rho,\tau)$, $p=(a_{1},a_{2},\dots,a_{k})$ with $t(a_{j})=o(a_{j+1})$ $(j=1,\dots,k-1)$, we define
@@
Set $\tilde{\rho}:=\rho\oplus\rho^{-1}$. A path $p=(a_{1},\dots,a_{s})$ with $a_{1}\in A_{br}$ and $a_{s}\in A_{is}$ in $G(\rho,\tau)$ can be represented by
$p=(e_{1},\Gamma_{1},e_{2},\Gamma_{2},\dots,e_{r},\Gamma_{r}),$
where $e_{1},\dots,e_{r}\in A_{br}$,
$\Gamma_{j}=(\tilde{\rho}^{0}(\xi_{j}),\tilde{\rho}^{1}(\xi_{j}),\dots,\tilde{\rho}^{k_{j}}(\xi_{j})),\;\;(\tilde{\rho}^{\ell}(\xi_{m})\in A_{is})$
with $e_{1}=a_{1}$ and $\rho^{k_{r}}(\xi_{r})=a_{s}$. Here if $\Gamma_{j}=\emptyset$, then $e_{j+1}=\bar{e}_{j}$. The inverse path of $p$ is defined by
$\bar{p}=(\bar{\Gamma}_{r},\bar{e}_{r},\dots,\bar{\Gamma}_{2},\bar{e}_{2},\bar{\Gamma}_{1},\bar{e}_{1}),$
where
$\bar{\Gamma}_{j}=(\tilde{\rho}^{k_{j}+1}(\xi_{j}),\tilde{\rho}^{k_{j}+2}(\xi_{j}),\dots,\tilde{\rho}^{d_{j}-1}(\xi_{j})).$
The inverse path of $p$ with $a_{1}\in A_{is}$ and $a_{s}\in A_{is}$ and so on is also defined in the same way.
@@
Set $\alpha,\beta\in\mathcal{A}$. If $\int_{c}\alpha=\int_{c}\beta$ for any closed path $c$ in $G(\rho,\tau)$, then $\int_{p}(\beta-\alpha)=\int_{p^{\prime}}(\beta-\alpha)$ for any path $p$ and $p^{\prime}$ with $o(p)=o(p^{\prime})$ and $t(p)=t(p^{\prime})$.
@@
It holds that $\int_{\bar{p}}\theta=-\int_{p}\theta$ for any path $p$ by the definition. Note that $p\cup\bar{p^{\prime}}$ is a closed path. Then we have
$0=\int_{p\cup\bar{p}}(\beta-\alpha)=\int_{p}(\beta-\alpha)+\int_{\bar{p^{\prime}}}(\beta-\alpha)=\int_{p}(\beta-\alpha)-\int_{p^{\prime}}(\beta-\alpha),$
@@
Lemma 5.1 tells us that the value $\int_{p}(\theta_{2}-\theta_{1})$ is independent of choice of path from $o(p)$ to $t(p)$. Let us fix a vertex $x_{*}$. Since $\int_{p}(\theta_{2}-\theta_{1})$ for a path $p$ starting from $x_{*}$ is determined by $t(p)=x$, we denote such a value by $\Delta(x)$, which is well-defined.
@@
for any $\epsilon\in A_{br}$, where $\xi^{\flat}=\mathrm{is}(\epsilon)$ and $\xi=\mathrm{is}^{\sharp}(\epsilon)$. Here $\theta_{1}$ is defined in (5.21). Now let us set $\theta_{2}$ by (5.22). Operation ()($\star$)( ⋆ ) keeps the parity of any paths passing through the vertex $x$, because $\tau(e)+\tau(e^{\prime})$ with $t(e)=x=o(e^{\prime})$ in $G$ is changed to
@@
Here for a path $p$ with $t(p)=x$ and the fixed vertex $x_{*}$,
$\Delta(x)=\int_{p}\theta_{2}-\theta_{1}$ (5.24)
is well-defined by Lemma 5.1. Inserting them into (5.23), we have
@@
The second equality derives from $o(\xi)=o(\epsilon)$ which gives the commutativity of the second diagonal matrix and $C$ in the first equality. Therefore introducing the unitary map $\mathcal{U}_{\Delta}$ by